### v2

# Understanding Ethereum Mempool Security under Asymmetric DoS by Symbolic Fuzzing

Abstract
In blockchains, mempool controls transaction flow before consensus, denial of whose service hurts the health and security of blockchain networks. This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool state space and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Compared to the baseline blockchain fuzzers, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits. Running mpfuzz on six major Ethereum clients leads to the discovering of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking. Rule-based mitigation schemes are proposed against newly discovered vulnerabilities.
1 Introduction
In Ethereum, a mempool buffers unconfirmed transactions from web3 users before they are included in the next blocks. Mempool provides the essential functionality to bridge the gap between varying rates of submitted transactions and rates of produced blocks, despite what types of transactions it serves, whether public or private transactions. As shown in recent studies [22], denying a mempool service can force the blockchain to produce blocks of low or even zero (Gas) utilization, undermining validators’ incentives and shrinking the blockchain networks in the long run, re-introducing the $51\%$ attacks. Besides, a denied mempool service can drop normal transactions, cutting millions of web3 users off the blockchain and failing the decentralized applications relying on real-time blockchain access.
Problem: Spamming the mempool to deny its service has been studied for long [17, 9, 27, 12]. Early designs by sending spam transactions at high fees burden attackers with high costs and are of limited practicality. What poses a real threat is Asymmetric DeniAl of Mempool Service, coined by ADAMS, in which the mempool service is denied at an asymmetrically low cost. That is, the attack costs in terms of the fees of adversarial transactions are significantly lower than those of normal transactions victimized by the denied mempool. In the existing literature, DETER [22] is the first ADAMS attack which works by sending invalid transactions to directly evict normal transactions in the mempool. MemPurge [29] is a similar mempool attack that finds a way to send overdraft transactions into Geth’s pending-transaction pool and cause eviction there. These known attacks are easy to detect (i.e., following the same direct-eviction pattern), and in fact, the DETER bugs reported in 2021 have been successfully fixed in all major Ethereum clients as of Fall 2023, including Geth, Nethermind, Erigon, and Besu. Given this state of affairs, we pose the following research question: Are there new ADAMS vulnerabilities in the latest Ethereum clients already patched against direct-eviction based attacks?
This work takes an automatic and systematic approach to discovering ADAMS vulnerabilities, unlike the existing DETER bugs that are manually found. Fuzzing mempool implementations is promising but also poses unique challenges: Unlike the consensus implementation that reads only valid confirmed transactions, the mempool residing pre-consensus needs to handle various unconfirmed transactions, imposing a much larger input space for the fuzzer. For instance, a mempool can receive invalid transactions under legitimate causes, [^1] For instance, future transactions can be caused by out-of-order information propagation in Ethereum. and factors such as fees are key in determining transaction admission outcomes. Existing blockchain fuzzers including Fluffy [30], Loki [23] and Tyr [19] all focus on fuzzing consensus implementation and don’t explore the extra transaction space required by mempool fuzzing. As a result, the naive application of a consensus fuzzer to fuzz mempool misses detecting the DETER bugs as evaluated in § 4.1, let alone discover more sophisticated new ADAMS bugs.
Proposed methods: To enable efficient mempool fuzzing, our key observation is that real-world mempool implementations admit transactions based on abstract “symbols”, such as pending and future transactions, instead of concrete value. Sending multiple concrete transactions generated under the same symbol triggers the same mempool behavior and should be avoided. Mutating one transaction under one symbol during fuzzing suffices to cover mempool behavior.
We propose symbolic stateful mempool fuzzing or mpfuzz. We first extract seven symbols from Ethereum semantics to cover future, parent, child, overdraft, latent overdraft, and replacement transactions. Then, mpfuzz generates a new transaction by symbol, sends the input to execute on a target mempool implementation, observes the end state of the mempool, and uses the symbolic state coverage as feedback to guide the next round of mutation. By this means, mpfuzz searches the symbolic mempool state space for the one matching the bug oracles induced from ADAMS. To further speed up the search in a limited time, we prioritize “promising” states by optimistically estimating the costs of unconfirmed transactions whose validity is subject to change in the future.
Found attacks: We run mpfuzz on six Ethereum execution-layer clients including Geth [2], Nethermind [5], Erigon [1], Besu [3], Reth [7], and OpenEthereum [6]. Compared to the baseline fuzzers, mpfuzz can find exploits faster by more than two orders of magnitude. On Ethereum clients of historical versions, mpfuzz can rediscover known DETER bugs. On the latest Ethereum clients where DETER bugs are fixed, mpfuzz can find new ADAMS attacks, described next.
We summarize the newly discovered ADAMS exploits into several patterns: 1) Indirect eviction by valid-turned-invalid transactions. Unlike the direct-eviction pattern used in DETER, the indirect-eviction attack works more stealthily in two steps: The attacker first sends normal-looking transactions to evict victim transactions from the mempool and then sends another set of transactions to turn the admitted normal-looking transactions into invalid ones, bringing down the cost. 2) Locking mempool, which is to occupy the mempool to decline subsequent victim transactions. Mempool locking does not need to evict existing transactions from the mempool as the eviction-based DETER attacks do. 3) Adaptive attack strategies where the attack composes adversarial transactions in an adaptive way to the specific policies and implementations of the mempool under test. Example strategies include composing transactions of multiple patterns in one attack, re-sending evicted transactions to evict “reversible” mempools, locking mempool patched against turning, etc.
We propose rule-based mitigation schemes against some of our discovered bugs and identify design challenges in mitigating others, which is left for future works.
All newly found ADAMS bugs are reported to the Ethereum Foundation with $7$ bugs confirmed and $3$ fixed in recent client releases [4]. Bug reporting is documented on the webpage [13] which also includes demos of mpfuzz on six Ethereum clients. mpfuzz will be open-sourced to facilitate vulnerability discovery in more and future clients.
Contributions: The paper makes contributions as follows:
$\bullet$ New fuzzing problem: This paper is the first to formulate the mempool-fuzzing problem to automatically find asymmetric DoS vulnerabilities as bug oracles. Fuzzing mempools poses new unique challenges that existing blockchain fuzzers don’t address and entails a larger search space including invalid transactions and varying fees.
$\bullet$ Symbolic fuzzing tool: The paper presents the design and implementation of mpfuzz, a symbolic stateful mempool fuzzer. Given a mempool implementation, mpfuzz searches the mempool state space using the feedback of symbolic state coverage and promising-ness to reach bug oracles. With a Python prototype and runs on real Ethereum clients, mpfuzz achieves a >100×>100$\times$> 100 × speedup in finding known DETER exploits compared to baselines.
$\bullet$ New mempool vulnerabilities: mpfuzz discovers new asymmetric-DoS vulnerabilities in six major Ethereum clients in the mainnet. By evaluation under real transaction workloads, all found attacks achieve $84.6-99.6\%$ success rates and low costs such as adversarial transaction fees 100×100$\times$100 × lower than the victim transaction fees.
2 Related Works
Consensus fuzzers: Fluffy [30] is a code-coverage based differential fuzzer to find consensus bugs in Ethereum Virtual Machines (EVM). Specifically, given an EVM input, that is, a transaction sequence, Fluffy sends it to multiple nodes running different Ethereum clients (e.g., Geth and OpenEthereum) for execution. 1) The test oracle in Fluffy is whether the EVM end states across these nodes are different, implying consensus failure. 2) Fluffy mutates ordered transaction sequences at two levels: it reorders/adds/deletes transactions in the sequence, and for each new transaction to generate, it randomly selects values in Gas limits and Ether amounts. For the data field, it employs semantic-aware strategies to add/delete/mutate bytecode instructions of pre-fixed smart contract templates. 3) Fluffy uses code coverage as feedback to guide the mutation.
Loki [23] is a stateful fuzzer to find consensus bugs causing crashes in blockchain network stacks. Specifically, Loki runs as a fuzzer node interacting with a tested blockchain node through sending and receiving network messages. 1) The test oracle in Loki is whether the tested blockchain node crashes. 2) The fuzzer node uses a learned model of blockchain network protocol to generate the next message of complying format, in which the content is mutated by randomly choosing integers and bit-flipping strings. In other words, the mutation in Loki is unaware of application level semantics. For instance, if the message is about propagating an Ethereum transaction, the transaction’s nonce, Ether amount, and Gas price are all randomly chosen; the sender, receiver, and data are bit flipped. 3) Loki records messages sent and received from the test node. It uses them as the state. New states receive positive feedback.
Tyr [19] is a property-based stateful fuzzer that finds consensus bugs causing the violation of safety, liveness, integrity, and other properties in blockchain network stacks. Specifically, in Tyr, a fuzzer node is connected to and executes a consensus protocol with multiple neighbor blockchain nodes. 1) The test oracle Tyr uses is the violation of consensus properties, including safety (e.g., invalid transactions cannot be confirmed), liveness (e.g., all valid transactions will be confirmed), integrity, etc. The checking is realized by matching an observed end state with a “correct” state defined and run by the Tyr fuzzer. 2) Tyr mutates transactions on generic attributes like randomly selecting senders and varying the amount of cryptocurrencies transferred (specifically, with two values, the full sender balance, and balance plus one). When applied to Ethereum, Tyr does not mutate Ethereum-specific attributes, including nonce or Gas price. It generates one type of invalid transaction, that is, the double-spending one, which in account-based blockchains like Ethereum are replacement transactions of the same nonce. 3) Tyr uses state divergence as the feedback, that is, the difference of end states on neighbor nodes receiving consensus messages in the same iteration. In Tyr, the states on a node include both the messages the node sent and received, as well as blocks, confirmed transactions, and other state information.
These existing blockchain fuzzers [19, 23, 30] cannot detect mempool DoS bugs and are different from this work. Briefly, mempool DoS does not trigger system crashes and can not be detected by Loki. Mempool content difference across nodes does not mean insecurity, and the mempool DoS bugs can not be detected by Fluffy (which detects consensus state difference). While Tyr aims to detect liveness and safety violations, their model of invalid transactions is rudimentary. Specifically, they only model double-spending transactions and cannot detect DETER bugs [22] that rely on more advanced invalid transactions, like future transactions and latent overdrafts.
Blockchain DoS: Blockchain DoS security has been examined at different system layers, including P2P networks [20, 24, 15, 28], mining-based consensus [25, 8], and application-level smart contracts [26, 18, 11] and DApp (decentralized application) services [21]. These DoSes are not related to mempools and are orthogonal to this work.
For mempool DoS, a basic attack is by sending spam transactions with high fees to evict victim transactions of normal fees [17, 9, 27, 12], which incur high costs to attackers. DETER [22] is the first asymmetric DoS on Ethereum mempool where the adversary sends invalid transactions of high fees (e.g., future transactions or latent overdrafts) to evict victim transactions of normal fees. Concurrent to this work is MemPurge [29] posted online in June 2023; in it, the attacker reconnects her adversarial future transactions and makes them latent-overdraft transactions. As we tested on the latest Ethereum clients in July 2023, such as Geth $v1.11.4$, the DETER bugs have been fixed. Both DETER and MemPurge are manually discovered.
3 Threat Model
In the threat model, an attacker aims at disabling the mempools on one or multiple nodes and causing the disruption of the transaction propagation service. This can lead to damage, such as delayed or dropped transaction inclusion and low-utilization blocks (even empty ones). The damage can benefit certain attackers in practice, such as competing RPC services, MEV bots, or block validators. Besides, the attacker aims at causing the above damage at asymmetrically low costs, that is, the fees of adversarial transactions included in the blockchain are lower than the fees of victim transactions outside the mempool, as will be presented.
The attacker has the same capability as in the DETER work [22], where the attacker joins an Ethereum network, discovers critical nodes (e.g., backend nodes of a RPC service, a node hosting MEV bot, or target block validators), and sends adversarial transactions.
3.1 Notations
Transactions: In Ethereum, we consider two types of accounts: a benign account of index $i$ denoted by $B_{i}$, and an adversarial account of index $i$ denoted by $A_{i}$. A (concrete) Ethereum transaction $tx$ is denoted by $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$}}$, where the sender is $s$, the nonce is $n$, the Gas price is $f$, and the transferred value in Ether is $v$. This work does not consider the “data” field in Ethereum transactions or smart contracts because smart contracts do not affect the validity of the Ethereum transaction. The notations are summarized in Table 1.
Table 1: Notations
Annotation
Meaning
Symbol
Meaning
$m$
Mempool length
$ops$
Tx sequence
$st$
State of residential txs in mempool
$dc$
History of declined txs from mempool
$A$
Adversarial tx sender
$B$
Benign tx sender
$tx\begin{bmatrix}{s}&{v}\\ {n}&{f}\end{bmatrix}$
A transaction from sender $s$, of nonce $n$, transferring Ether value $v$, and with Gas Price $f$
We use each of the following symbols to represent a disjoint group of transactions. Informally, $\mathcal{N}$ is a transaction sent from a benign account, $\mathcal{F}$ a future transaction, $\mathcal{P}$ a parent transaction, $\mathcal{C}$ a child transaction, $\mathcal{O}$ an overdraft transaction, $\mathcal{L}$ a latent overdraft transaction, and $\mathcal{R}$ a replacement transaction. These symbols are formally specified in § 4.
Particularly, we use the notion of latent overdraft from the existing work [22], which indicates a child transaction that by itself does not overdraft but does overdraft when taking into account its parent transactions. Suppose Alice has a balance of $5$ Ether and sends $tx_{1}$ of nonce $1$ spending $2$ Ether and $tx_{2}$ of nonce $2$ spending $4$ Ether. $tx_{2}$ is a latent overdraft.
States: We recognize two mempool-related states in Ethereum, the collection of transactions stored in mempool $st$, and the collection of declined transactions $dc$.
Suppose a mempool of state $\langle{}st_{i},dc_{i}\rangle{}$ receives an arriving transaction $tx_{i}$ and transitions to state $\langle{}st_{i+1},dc_{i+1}\rangle{}$.
An arriving transaction can be admitted with eviction, admitted without eviction, or declined by a mempool. The three operations are defined as follows. 1) An arriving transaction $tx_{i}$ is admitted with evicting a transaction $tx_{i}\textquoteright$ ’ from the mempool, if $st_{i+1}=st_{i}\setminus{}tx_{i}^{\prime}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 2) $tx_{i}$ is admitted without evicting any transaction in the mempool, if $st_{i+1}=st_{i}\cup{}tx_{i},dc_{i+1}=dc_{i}$. 3) $tx_{i}$ is declined and does not enter the mempool, if $st_{i+1}=st_{i},dc_{i+1}=dc_{i}\cup{}tx_{i}$.
3.2 Bug Oracles
Suppose a mempool of initial state $\langle{}st_{0},dc_{0}\rangle{}$ receives a sequence of transactions, denoted by $ops$. Here, we simplify the problem by not considering arriving blocks during fuzzing. We define two bug oracles: eviction-based attacks and declining-based locking attacks.
Definition 3.1 (Eviction bug oracle)
A mempool execution from initial state $st_{0}$ to end state $st_{n}$ under transaction sequence $ops$ is a successful ADAMS eviction attack, if.f. 1) the attack damage is full, that is, end state $st_{n}$ contains only adversarial transactions and initial state $st_{0}$ contains only benign transactions, and 2) the attack cost is low, that is, the total adversarial transaction fees in the end states are smaller, by a multiplicative factor $\epsilon$, than the attack damage measured by the victim transaction fees in the initial state. Formally,
$st_{0}\cap{}(st_{n})$ $=$ $\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\epsilon\sum_{tx\in{}st_{0}}tx.fee$ (1)
Definition 3.2 (Locking bug oracle)
A mempool execution from initial state $\langle{}st_{0},dc_{0}\rangle{}$ to end state $\langle{}st_{n},dc_{n}\rangle{}$ under transaction sequence $ops$ is a successful ADAMS locking attack if.f. 1) the attack damage is full, that is, the victim normal transactions in $ops$ are completely declined by the mempool storing adversarial transactions, where normal transactions and adversarial transactions are sent by different accounts, and 2) the attack cost is low, that is, the total adversarial transaction fees in the mempool are smaller, by a multiplicative factor $\lambda$, than the attack damage measured by the fees of victim transactions declined. Formally,
$\cup_{txd\in{}dc_{n}}{txd.sender}$$\cap$ $\cup_{tx\in{}s_{n}\cup{}ops\setminus{}dc_{n}}{tx.sender}=\varnothing$
$\sum_{tx\in{}st_{n}}tx.fee$ $<$ $\lambda\sum_{tx\in{}dc_{n}}tx.fee$ (2)
4 Stateful Mempool Fuzzing by mpfuzz
4.1 Motivating Observations
(In)feasibility of code-coverage-only stateless fuzzing: To fuzz a mempool, a baseline design is to use a code-coverage-based fuzzer. In each iteration, the fuzzer generates a bit string, parses it into a sequence of transactions and sends them to the tested mempool for execution. Upon each end state, it checks whether the test oracle defined above is satisfied. This design takes code coverage as the only feedback and is stateless because the feedback does not consider the end state.
The stateless design would be ineffective in fuzzing mempool or finding ADAMS exploits. For example, consider fuzzing a three-slot mempool to find a DETER-X exploit consisting of three future transactions [22]. Suppose the current mempool input is a sequence of one future transaction. After mutation, it may try an input of two future transactions, which, however, does not increase code coverage (the second future transaction is executed by the same code path in a mempool as the first future transaction). Thus, the fuzzer discards the two-future-transaction input and misses finding the three-future-transaction exploit. We validate the ineffectiveness of stateless fuzzer by experiments in § 6.
A more promising design is stateful mempool fuzzing, in which the mempool state is included in the feedback to guide next-iteration fuzzing. In the previous example of finding a three-future-transaction exploit (DETER-X), sending two future transactions leads to a different mempool state than sending one. A simple state-coverage-based fuzzer would view this as positive feedback and further explore this direction toward finding the three-future-transaction exploit.
(In)feasibility of existing consensus fuzzers: Existing consensus fuzzers [19, 23, 30] cannot detect mempool DoS or ADAMS bugs. Specifically, Loki’s test oracle is detecting system crashes, while a successful ADAMS attack does not necessarily trigger system crashes. Using Loki to detect ADAMS would cause many false negatives.
Fluffy detects the unsynchronizable difference of post-consensus states as vulnerability. However, the mempools in two clients, say Geth and Besu under benign transactions, would be different. In other words, post-consensus state differences across clients could indicate insecurity, but the difference in pre-consensus states across clients, like mempool, does not mean insecurity. Using Fluffly to detect ADAMS could cause false positives.
Tyr’s test oracle detects property violations on post-consensus states. Particularly, while Tyr models state liveness, their definition “valid transactions should be executed eventually” is, unfortunately, over-simplifying and fails to model the legitimate pre-consensus cases in real Ethereum clients; for instance, valid transactions of low Gas price can be dropped by Ethereum mempools and will not be eventually executed. Besides, they model double-spending transactions as the only type of invalid transaction. In Ethereum, double-spending transactions are of the same nonce. And there are many more sophisticated invalid transactions in Ethereum that Tyr does not model, including future transactions, latent overdrafts, etc.
We analyze the case of existing consensus fuzzers in detecting known DETER attacks [22]. Loki cannot detect DETER because DETER does not cause system crashes. Fluffy does not model future transactions and can not find the DETER attacks that rely on future transactions. Similarly, Tyr that does not model sophisticated invalid transactions would miss detecting DETER attacks.
Challenges: In stateful mempool fuzzing, the essential design problem is how to generate the next input transaction given a current “state”. State $st$ is the collection of transactions in the mempool and those included in the blockchain.
The key challenge is the large input space of transactions. Recall that an Ethereum transaction consists of multiple attributes (sender, nonce, Gas price, value, etc.), each defined in a large domain (e.g., 64-bit string). Exploring the raw transaction space is hard; the comprehensive search is inefficient, and randomly trying transactions as done in the state-of-the-art blockchain fuzzers (in § 2) is ineffective.
4.2 Proposed Fuzzer Design
4.2.1 Transaction Symbolization
Intuition: We propose symbolizing transactions for efficient and effective mempool stateful fuzzing. Our key idea is to map each group of concrete transactions triggering an equivalent mempool behavior into a distinct symbol so that searching one transaction is sufficient to cover all other transactions under the same symbol. Transaction symbolization is expected to reduce the search space from possible concrete transactions to the symbol space.
Table 2: Symbols, transactions and associated costs. Tx refers to the instantiated transaction under a given symbol.
Symbol Description
Tx
$cost()$ $opcost()$
$\mathcal{N}$ Benign
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$}}$
$3$ $3$
$\mathcal{F}$ Future
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{P}$ Parent
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A>r}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$
$[4,m]$ $[4,m]$
$\mathcal{C}$ Child
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
$m+1$ $1$
$\mathcal{O}$ Overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{L}$ Latent overdraft
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$
00 00
$\mathcal{R}$ Replacement
$tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$
00 00
Specification: Suppose the current state contains adversarial transactions sent from $r$ accounts $A_{1},\dots,A_{r}$. These accounts have an initial balance of $m$ Ether, where $m$ equals the mempool capacity (say $m=1000$).
Symbol $\mathcal{N}$ defines a transaction subspace covering any normal transactions sent from any benign account and of any nonce, namely $tx\begin{bmatrix}{B*}&{*}\\ {*}&{*}\end{bmatrix}$ In mpfuzz, Symbol $\mathcal{N}$ is instantiated to concrete transactions of fixed Gas price $3$ wei and of value $1$ Ether. That is, symbol $\mathcal{N}$ is instantiated to transaction $tx\begin{bmatrix}{B*}&{1}\\ {*}&{3}\end{bmatrix}$ , as shown in Table 2. Among all symbols, $\mathcal{N}$ is the only symbol associated with benign sender accounts.
Symbol $\mathcal{F}$ defines the transaction subspace of any future transaction sent from an adversarial account; the future transaction is defined w.r.t. the current state. mpfuzz instantiates symbol $\mathcal{F}$ to concrete transactions of nonce $m+1$, value $1$ Ether, Gas price $m+1$ wei, and any adversarial account $*A*$. The instantiating pattern for symbol $\mathcal{F}$ is $tx\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$.
Symbol $\mathcal{P}$ defines the transaction subspace of any parent transaction from an adversarial account w.r.t. the current state. mpfuzz instantiates $\mathcal{P}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A\geq{r}}&{1}\\ {1}&{[4,m]}\end{bmatrix}$}}$. Here, the transaction is fixed with a Gas price in the range $[4,m]$ wei, value at $1$ Ether, and nonce at $1$.
Symbol $\mathcal{C}$ defines the transaction subspace of any adversarial child transaction w.r.t. the state. mpfuzz instantiates $\mathcal{C}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. The instantiated child transaction is fixed at Gas price of $m+1$
Symbol $\mathcal{O}$ defines the transaction subspace of any adversarial overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{O}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m+1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. Recall that all adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{L}$ defines the transaction subspace of any adversarial latent-overdraft transaction w.r.t. state $st$. mpfuzz instantiates $\mathcal{L}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {\geq{}2}&{m+1}\end{bmatrix}$}}$. All adversarial accounts have an initial balance of $m$ Ether.
Symbol $\mathcal{R}$ defines the transaction subspace of any adversarial “replacement” transaction w.r.t., the state. Unlike the symbols above, the transactions under Symbol $\mathcal{R}$ must share the same sender and nonce with an existing transaction in the current state $st$. mpfuzz instantiates $\mathcal{R}$ to a transaction of Pattern $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A[1,r]}&{m-1}\\ {1}&{m+1}\end{bmatrix}$}}$. Here, we simplify the problem and only consider replacing the transaction with nonce $1$.
Figure 1: Symbols and transaction space reduction.
Given that a concrete transaction in this work consists of four attributes, transactions are defined in a four-dimensional space. We visualize transaction symbols (with an incomplete view) in two two-dimensional spaces in Figure 1, that is, one of transaction sender and nonce and the other of sender and value. For each symbol, the figures depict the transaction subspace covered by the symbol (in white shapes of black lines) and the transaction pattern instantiated by mpfuzz under the symbol (in red lines or shapes). In our design, the instantiated transactions are a much smaller subset of the defining transaction space. For instance, given state $st$ of transactions sent from account $A_{1},\dots,A_{r}$, transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ is within the defining space of symbol $\mathcal{F}$, but mpfuzz does not instantiate $\mathcal{F}$ by transaction $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{r+1}}&{*}\\ {3}&{*}\end{bmatrix}$}}$ (but instead to transactions $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A*}&{1}\\ {m+1}&{m+1}\end{bmatrix}$}}$).
4.2.2 State Search Algorithm
Algorithm 1 mpfuzz(SeedCorpus $sdb$, Mempool $mp$)
1:$sdb$.init($mp$);
2:while !$sdb$.is_empty() or timeout do
3:    $st,ops$ = $sdb$.next(); $\triangleright$ Selection by energy
4:    for all $ops^{\prime},st^{\prime}$ = mutateExec($ops,st$do
5:         if is_ADAMS($st^{\prime}$then $\triangleright$ Test oracle
6:             emit($ops^{\prime}$, $st^{\prime}$, "Found an exploit");
7:         else
8:             if !$sdb$.feedback($st^{\prime}$, $sdb$then
9:                 $sdb$.add($st^{\prime}$, $ops^{\prime}$);
10:             end if
11:         end if
12:    end for
13:end while
Algorithm overview: We propose a stateful fuzzing algorithm, listed in Algorithm 1, to selectively explore the input space in a way that prioritizes new and promising mempool states towards triggering ADAMS oracle.
The core data structure is a seed corpus or $sdb$ which stores a list of input-state pairs or seeds. Upon running the algorithm, the corpus maintains the mempool states covered so far by the algorithm execution. The algorithm runs an outer loop that continues until the corpus is empty, or only the states of zero energy are left, or timeout. In each iteration, the algorithm retrieves the next seed from the corpus based on how promising the seed is in reaching a state triggering the oracle (a.k.a., the energy [16]). A seed consists of a symbolic input $ops$, which consists of transaction symbols, and a symbolic state $st$ reached by running a transaction sequence instantiated from the input $ops$ (Line 3). The algorithm runs an inner loop, in each iteration of which it mutates the input in the current seed ($\langle{}st,in\rangle{}$) and executes the mutated input against a reinitialized empty mempool, producing end state $st^{\prime}$ (Line 4). If the end state satisfies ADAMS conditions (i.e., the test oracle), it emits the mutated input-state pair $ops^{\prime},st^{\prime}$ as a newly found exploit. The algorithm further checks the feedback: The feedback is positive if the mutated input $ops^{\prime}$ brings the mempool state $st^{\prime}$ closer to triggering the test oracle than $st$. This entails mutated state $st^{\prime}$ to be different from state $st$ (i.e., increased state coverage) and state $st^{\prime}$ to achieve larger damage or lower attack cost. In case of positive feedback, the algorithm would add mutated input-state pair $ops^{\prime},st^{\prime}$ to the corpus.
This algorithm assumes the mempool is deterministic. That is, given a seed $\langle{}ops,st\rangle{}$, running the same input $ops$ against an empty mempool multiple times always results in the same end state $st$. In practice, we generate inputs to avoid the non-deterministic behavior of real mempool implementations.
The algorithm can be configured with initial seeds and input-mutation strategies. By default, we use one initial seed whose input fills up the mempool with normal transactions. The default input-mutation strategy is to append the current input with a newly generated transaction.
Algorithm 2 mutateExec(SymbolInput $ops$, SymbolState $st$)
1:$ops^{\prime}$=mutateSymbol($ops$);
2:$st_{c}$=executeUnappended($ops,ops^{\prime}$);
3:$st_{c}^{\prime}$=executeAppended($ops^{\prime},st_{c},st$);
4:return $st_{c}^{\prime}$
Input mutation: Given a symbolized input $ops$ and symbolized state $st$, the mutation algorithm is to explore each “slightly” different input $ops^{\prime}$ and its associated state $st^{\prime}$, such that the next stage can find the inputs producing positive feedback.
Internally, the algorithm proceeds at two levels, that is, symbols and concrete value. Specifically, as shown in Algorithm 2, 1) it appends the symbolized input $ops$ with a previously untried symbol, generating the “mutated” symbolized input $ops^{\prime}$ (Line 1 in Algorithm 2). In this step, the algorithm tries different symbols in the following order: $\mathcal{P},\mathcal{L},\mathcal{C}$. 2) It instantiates symbolized $ops^{\prime}$ to a transaction sequence and executes the sequence in the tested mempool to obtain the concrete end state (Line 2 in Algorithm 2). This step is necessary for instantiating and executing the mutation in the next step. Specifically, in this step, the algorithm instantiates symbolized input $ops$ to a concrete input $in_{c}$, which is a sequence of transactions. It then drives the transactions to a reinitialized mempool for execution. It returns the resulting concrete state $st_{c}$. 3) The algorithm instantiates the appended symbols under the context of the previous state $st$ and $st_{c}$. At last, it executes the appended transaction on mempool $st_{c}$ to obtain the concrete end state $st^{\prime}_{c}$ under the mutated input (Line 3 in Algorithm 2).
Mutation feedback: We describe how mempool states are symbolized before presenting state-based feedback. In mpfuzz, a symbolized state $st$ is a list of transaction symbols, ordered first partially as follows: $\mathcal{N}\preceq{}\mathcal{E}\preceq{}\mathcal{F}\preceq{}\{\mathcal{P},\mathcal{L},\mathcal{C}\}$. $\mathcal{E}$ refers to an empty slot. State slots of the same symbols are independently instantiated into transactions, except for $\mathcal{P},\mathcal{C},\mathcal{L}$. Child transactions (i.e., $\mathcal{C},\mathcal{L}$) are appended to their parent transaction of the same sender (i.e., $\mathcal{P}$). Across different senders, symbols are ordered by the parent’s GasPrice. For instance, suppose a mempool stores four concrete transactions, $tx\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$, $tx\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$, $tx\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$, $tx\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$ , they are mapped to four symbols, $\mathcal{P},\mathcal{F},\mathcal{P},\mathcal{C}$, and are further ordered in a symbolic state by $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$.
In mpfuzz, the feedback of an input $ops^{\prime}$ is based on the symbolic end state $st^{\prime}$. Positive feedback on state $st^{\prime}$ is determined conjunctively by two metrics: 1) State coverage that indicates state $st^{\prime}$ is not covered in the corpus, and 2) state promising-ness that indicates how promising the current state is to reach a state satisfying bug oracles.
$feedback(st^{\prime},st,sdb)=1$ , if $\text{if }st\_coverage(st^{\prime},sdb)=1\land{}$ (3)
$st\_promising(st^{\prime},st)=1$
Specifically, state coverage is determined by straightforwardly comparing symbolic state $st^{\prime}$, as an ordered list of symbols, with all the symbolic states in the corpus. For instance, state $\mathcal{F}\mathcal{P}\mathcal{C}\mathcal{P}$ is different from $\mathcal{F}\mathcal{P}\mathcal{P}\mathcal{C}$. This also implies concrete transaction sequence $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{5}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10001}\end{bmatrix}$}}$ is the same with $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {1}&{4}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{B}&{*}\\ {100}&{100}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{2}}&{*}\\ {1}&{6}\end{bmatrix}$}}$, $tx{\hbox{\scriptsize$\displaystyle\begin{bmatrix}{A_{1}}&{*}\\ {2}&{10}\end{bmatrix}$}}$, as they are both mapped to the same symbolic state.
How promising a state $st^{\prime}$ is (i.e., $st\_promising(st^{\prime})$) is determined as follows: A mutated state $st^{\prime}$ is more promising than an unmutated state $st$ if any one of the three conditions is met: 1) State $st^{\prime}$ stores fewer normal transactions (under symbol $\mathcal{N}$) than state $st$, implying more transactions evicted and higher damage (i.e., $evict\_normal(st^{\prime},st)=1$). 2) State $st^{\prime}$ declines (speculatively) more incoming normal transactions than state $st$, also implying higher damage (i.e., $decline\_normal(st^{\prime},st)=1$). 3) The total fees of adversarial transactions in state $st^{\prime}$ are lower than those in state $st$, implying lower attack costs. We consider two forms of costs: concrete cost and symbolic cost. The former, denoted by $cost(st^{\prime})$, is simply the total fees of adversarial transactions instantiated from a symbolic state $st^{\prime}$. The latter, denoted by $opcost(st^{\prime})$, optimistically estimates the cost of transactions in the current state $st^{\prime}$ that contributes to a future state triggering test oracle. We will describe the symbolic cost next. Formally, state promising-ness is calculated by Equation 4.
$st\_promising(st^{\prime},st)=$ $1,$ if $\text{if }evict\_normal(st^{\prime},st)=1\lor{}$ (4)
$decline\_normal(st^{\prime},st)=1\lor{}$
$cost(st^{\prime})<cost(st)\lor{}$
$opcost(st^{\prime})<opcost(st)$
Seed energy: Recall that in mpfuzz, the next seed is selected from the corpus based on energy. By intuition, the energy of a state is determined based on how promising the state is in triggering the test oracle. In addition to the state promising-ness used in deciding feedback, state energy incorporates fuzz runtime information, such as how many times the state has been selected.
Specifically, the energy of a seed $\langle{}ops,st\rangle{}$ is determined by Equation 5.
$Energy(in,st)$ $=$ $b/opcost(st)$ (5)
First, each seed in the corpus records how many times it has been mutated. The more symbols it has mutated in the past, the less energy the seed currently has and the lower priority it will be selected next. $b$ can be configured differently to traverse the state tree differently. In particular, breadth-first search (BFS) is by the following configuration: $b=1$ if at least one symbol has not been tried (for mutation) in the current seed. Otherwise, $b=0$.
Second, we use the symbolic cost to estimate how promising a state is and use it in seed energy.
Estimate state cost: When mpfuzz determines how promising a state is, it needs to look beyond the current state and into all possible descendant states. We propose a heuristic that optimistically estimates the descendant-state costs given a current state. The key intuition is the following: In Ethereum, the validity of a child transaction (under symbol $\mathcal{C}$) depends on its parent/ancestor transaction. Thus, even though a transaction in the current state is valid, the transaction can be “turned” into an invalid one in subsequent states. We thus attribute, optimistically, the cost of transaction $\mathcal{C}$ to value $1$, so that it is preferable to a parent transaction $\mathcal{P}$ in input mutation and seed selection, and turning $\mathcal{C}$ into an invalid one like $\mathcal{L}$ or $\mathcal{F}$ is also encouraged. The cost profiles used in mpfuzz are summarized in Table 2.
5 Found Exploits
Running mpfuzz on six leading Ethereum execution-layer clients, namely, Geth, Besu, Nethermind, Erigon, Reth, and OpenEthereum, leads to the discovery of $22$ bugs on all six Ethereum clients, including $7$ conforming to the known DETER attacks on clients of historical versions and $15$ new bugs on the clients of the latest versions. We describe two selected patterns of these bugs as follows.
5.1 Found Eviction Attacks (Selected)
Exploit $XT_{4}$: Indirect eviction by valid-turned-overdraft transactions: This eviction attack works in two steps: 1) The attacker first sends valid transactions at high fees to evict normal transactions initially stored in the mempool. These transactions are sent from $k$ accounts, each of which sends $l$ transactions. 2) She then sends $k$ transactions; each of the transactions is of nonce $1$, at a high fee, of high value $v_{1}$, and from the same $k$ accounts in Step 1). The transactions would replace the transaction of the same sender and nonce sent in Step 1). Once the replacement is finished, they turn their child transactions into latent overdraft. Specifically, if a sender’s balance is $bal$ and the Ether value of transaction of nonce 2 is $v_{2}$, then $v_{1}$ is carefully crafted to enable turned latent overdraft, that is, $v_{1}<bal$ and $v_{1}+v_{2}>bal$.
In practice, many clients are found vulnerable under this exploit, including Geth $<v1.11.4$, Erigon, Besu, Nethermind, and OpenEthereum.
Exploit $XT_{6}$: Compositional indirect eviction (by multi-round valid-turned-future transactions): The latest Geth $\geq{}v1.11.4$ is patched against direct eviction attacks. To evade the protection, this eviction attack uses indirect eviction.
The attack works in two steps: 1) It first control-evicts the mempool using valid-turned-future transactions to avoid the triggering condition of its limit of transactions per sender. 2) It then conducts another run of valid-turned-future transactions, this time with transactions sent from one account, namely $1\times{}l$. It can succeed since the limit is not triggered. The overall attack cost is the fee of one transaction; compared to the single run of valid-turned-future attack, $XT_{6}$’s cost is much lower.
5.2 Found Locking Attacks (Selected)
[…]
While a FIFQ mempool can not be vulnerable to eviction attacks, it can be easily locked by Exploit $XT_{8}$ as follows: Whenever the attacker observes an empty slot present in the mempool, she sends a pending transaction of a minimally necessary fee to occupy the slot. Any transactions that arrive after will encounter a full mempool and are declined.
6 Performance Evaluation of mpfuzz
Preliminary of GoFuzz: Upon each fuzzing iteration, GoFuzz [10] generates a bit-string of variable length, feeds it to our fuzzer code, and receives from our fuzzer a binary value of feedback, be either positive or negative. The GoFuzz further checks if this run increases the code coverage in the Geth client (as will be described, our fuzzer invokes functions in Geth’s txpool). If the feedback is positive and code coverage increases, GoFuzz adds the current bit-string to the seeds before running into the next iteration.
Baseline B1: Stateless fuzzer. We implement a stateless fuzzer in the GoFuzz framework described above. 1) Given a bit string generated by GoFuzz, our fuzzer code parses it into a sequence of transactions; the length of the sequence depends on the length of the bit string. It skips a bitstring that is too short. 2) The fuzzer then sets up and initializes a Geth txpool instance with $m$ normal transactions, each sent from a distinct account, of nonce $1$, of value $1$ Ether (lower than the account balance), and of Gas price $3$ wei. 3) The fuzzer sends the transaction sequence to the initialized mempool for execution. 4) It checks the mempool state after execution: If the eviction-based test oracle (Equation 1) is met, it emits the current transaction sequence as an exploit.
Baseline B2: Concrete-state-coverage fuzzer. We also build the second baseline, a mempool fuzzer that takes concrete state coverage as feedback. Specifically, in each iteration, the fuzzer appends to the current transaction sequence a new transaction. Given the $m$-slot mempool, the fuzzer tries $m$ values for senders, $m$ values for nonces, $m$ values for Gas price, and $m$ values for Ether amount. After sending to the mempool the current transaction sequence, it sorts the transactions in the mempool by senders and nonces. It then hashes the sorted transactions in the mempool and checks the presence of the hash digest in the seeds. If no hash exists, the current transaction sequence would increase the concrete-state coverage and is inserted to the corpus. The corpus is a FIFO queue where the seed inserted earlier has high priority to be de-queued.
Baseline B3: Invalid-tx energy fuzzer. We build the last baseline that is similar to the concrete-state-based fuzzer except for the following: It uses as the energy the number of invalid transactions on a mempool state. The seed with more invalid transactions in the mempool has higher energy or priority to be selected for the next round of fuzzing.
We implement Baseline B1 in Go/GoFuzz and implement B2 and B3 in Python. We use a test oracle that can detect DETER attacks, that is, the number of invalid transactions in a mempool state equal $m-1$.
We run the three baselines and mpfuzz against a Geth client under two settings: A small setting where the mempool is resized to $6$ slots $m=6$ and all fuzzers run for two hours (if test oracle is not triggered), and a medium setting where $m=16$ and fuzzers run for $16$ hours.
Table 3: Fuzzing Geth $v1.11.3$’s mempool by different approaches to detect ADAMS exploit.
Settings
B1 B2 B3 mpfuzz
6-slot-2h
timeout $54$ (min) $8$ (min)
$0.11$ (min)
16-slot-16h
timeout timeout $447$ (min)
$0.84$ (min)
Table 3 reports the results under the small and medium settings. For the small setting with $6$ slots, the stateless fuzzer (baseline B1) cannot find any DETER attack in two hours, concrete-state-coverage fuzzer (baseline B2) can find a DETER exploit in 54 minutes, and the stateful fuzzer based on invalid-transaction counter finds a DETER exploit in $8$ minutes. By contrast, mpfuzz finds a DETER exploit in $0.11$ minutes.
For the medium setting, with $6$ slots, both baseline B1 and B2 cannot find any DETER exploit in $16$ hours, and baseline B3 can find a DETER exploit in $7.4$ hours. By contrast, mpfuzz finds a DETER exploit in under a minute.
We further run our mpfuzz on different clients, including Geth $v1.10.11$, Geth $v1.11.4$, Erigon $v2.42.0$, and Nethermind $v1.18.0$. The tested mempool is configured at $16$ slots. We report the number of exploits found by mpfuzz in a $16$-hour period. Figure 3 presents the result that mpfuzz finds $4096$ exploits in the first $8$ hours on Geth-$v1.10.11$. On Geth-$v1.11.4$, mpfuzz quickly finds one exploit (i.e., exploit $XT_{6}$ as described in § 5.1) within $2$ minutes and does not find anymore in the next $16$ hours. On Nethermind, it finds one exploit (exploit $XT_{1}$) within two minutes and finds the next one (exploit $XT_{4}$) near the end of $16$-th hour. On Erigon, it finds one exploit ($XT_{4}$) in the $16$-th hour.
Figure 2: # exploits found
Figure 2: # exploits found
Figure 3: # states covered
Figure 3 shows the number of states that are explored by mpfuzz on a $16$-slot mempool in 16 hours. mpfuzz explores $749$ and $91428$ states on Geth $v1.10.11$ in $2$ minutes and $16$ hours respectively. On Geth $v1.11.4$, mpfuzz explores $372$ and $23286$ states in $2$ minutes and $16$ hours respectively. However, the performance of Erigon and Nethermind is much lower than that of Geth. mpfuzz explores $12$ and $1239$ states on Erigon in $2$ minutes and $16$ hours respectively. On Nethermind, mpfuzz explores $15$ and $1433$ states in $2$ minutes and $16$ hours, respectively.
The reason mpfuzz is more performant on Geth is that we implemented an external API on Geth to initialize the mempool; in each iteration of fuzzing on Geth, mpfuzz calls the API to initialize the mempool. In contrast, on Nethermind and Erigon, mpfuzz restarts the client in each iteration, which is consuming.
7 Evaluation of Discovered Attacks
This section shows the evaluation of turning attacks and locking attacks.
7.1 Evaluation of Turning Attacks
We design experiments to measure the success of ADAMS attacks in real Ethereum networks.
7.1.1 Evaluation in Testnet
Setup: We set up our experiment platform by connecting the attacker node to the Goerli testnet. The attacker node runs an instrumented Geth client at the execution layer which sends crafted transactions to the testnet.
To monitor transaction propagation, we launch an independent supernode (called measurement node) in the same testnet. The measurement node runs a reconfigured Geth client where the limit of peers/neighbors is removed and the measurement node can connect to as many neighbors as possible, hence a super node. The measurement node is not (directly) connected to the attacker node. When setting up our experiments, we ran the measurement node in Goerli for seven days and found the node is stabilized at $290$ neighbors.
(a) Transaction propagation in six messages
(c) Etherscan screenshot of the blocks generated during the attack on Goerli (txs in six messages)
Figure 4: Mounting $XT_{4}$ attacks on Goerli: Emptying blocks and propagating malicious transactions to the entire network.
Experiment results: The instrumented measurement node is able to log the received messages from different neighbors; these messages include those of transactions and of transaction hashes. The measurement node could receive the same transaction from different neighbors, and the log stores the transaction-neighbor pairs. We started logging the received messages one month before the experiment.
To do an experiment, we make the attacker node send $XT_{4}$ transactions using $384$ accounts. Each account sends $16$ valid pending transactions, followed by a replacement transaction. In total, the attacker node sends $6144$ valid transactions and $384$ replacement transactions. The Gas prices of the valid and replacement transactions are set to be $8$ Gwei and $130$ Gwei, respectively. In one setting, we split the $6528$ transactions into six messages (i.e., the first five being $1229$ transactions each and the last one with $384$ transactions). We send these messages in order and wait for $1.5$ seconds between two consecutive messages. For comparison, in another setting, we send the $6144$ valid transactions in one message and the $384$ replacement transactions in another message. Finally, we wrap up the experiment by waiting after the attacker node sends all messages and all replacement transactions are included in the blockchain. We then turn off the measurement node’s logging and search for the presence of all $6528$ transactions across neighbor nodes in the log. Given a transaction nonce, we report the percentage of neighbor nodes in the testnet that have propagated it to the measurement node.
Table 4: Types of transactions included in the four blocks during the attack in Figure 4c. ‘# private tx’ means the number of private transactions.
# block
# private txs
# public txs
# attack txs
$8040010$
$5$
00
$384$
$8040009$
$78$
$4$
$384$
$8040008$
$1$
00
$384$
$8040007$
$5$
00
$384$
We run experiments of six-message and two-message propagation, repeat each experiment by $3$ times, and report the average result of node percentage in Figure 4a and Figure 4b. Figure 4a shows the transaction propagation with six messages. More than 90% of the testnet nodes have received all of the $15$ valid transactions in a sequence and the replacement transactions. The only exception is the relatively low percentage for the $16$-th transaction of the maximal nonce. By comparison, when sending the transactions in two messages, the propagation fails as shown in Figure 4b where nodes did not propagate any transaction of nonce larger than $5$.
Additionally, Figure 4c shows the generated blocks in the experiment propagating six messages. We took the screenshot from etherscan.io and labeled it in red with information regarding the attack. Before the attack begins, the Goerli testnet normally utilizes $24-47\%$ of the Gas in a block. For ethical consideration, we limit our attack to be short, specifically, four blocks long.
Right after the launch of the attack, in block $8040007$, the Gas used by normal transactions drops to $3.19\%$ and the Gas used by adversarial transactions (denoted by red bars in Figure 4c) is $26.88\%$. The included adversarial transactions are $384$ replacement transactions sent in the second round of $XT_{4}$. The $6144-384=5760$ child transactions sent in the first round are not included in the block. Other blocks during the attack, namely $8040008$ and $8040010$, are similar. In the third block $8040009$, $73.1\%$ Gas is spent on including $82$ normal transactions as shown in Table 4 not sent by our attacker node. To explain it, we inspect the raw history of transactions arriving at and logged by our measurement supernode and found that out of the $82$ included normal transactions, only $4$ are present in the log, implying the other $78$ normal transactions are private ones that were not broadcast to the measurement node. Furthermore, the private mempool in today’s Ethereum clients [14] has a capability of no limit, which, while being feasible under today’s low private-transaction workloads, can cause resource-exhaustion vulnerability. Notice that we did not discover specific critical nodes to attack as needed by DETER attacks [22].
8 Discussions
[…]
Defenses: Rule-based mitigation can fix some of our discovered bugs: $XT_{4}$ is fixed by detecting latent overdraft transactions upon replacement. Locking-based $XT_{8}$ is fixed by re-enabling eviction and replacement; to avoid reintroducing eviction vulnerabilities, the fix needs to comply with the security rules against eviction attacks.
Ethical concerns: When evaluating attacks, we only mounted attacks on the testnet and did so with minimal impacts on the tested network. For instance, our attack lasts a short period of time, say no more than $4$ blocks produced. We did not test our attack on the Ethereum mainnet. When reporting bugs, we disclose to the developers the mitigation design tradeoff and the risk of fixing one attacks by enabling other attacks (described above). To prevent introducing new bugs, we did not suggest fixes against locking ($XT_{8}$).
Private and public mempools: In Ethereum, transactions can be sent privately or publicly. A private transaction received on a block builder is directly packed to a block without broadcast to other builders. By contrast, a public transaction received on a node is broadcast to other nodes before it is included in the next block.
On the one hand, whether a mempool stores private or public transactions is orthogonal to the admission/eviction policy the mempool runs. Empirically, we found three open-source Ethereum block builders, named EigenPhi builder, Flashbot builder $v1.11.5-0.2.1$ and bloXroute builder. Specifically, Flashbot and bloXrout builder run the same admission policy as Geth $v1.11.4$, while EigenPhi builder runs the same admission policy as Geth $<v1.11.4$.Therefore, these private mempools are subject to the same vulnerabilities as the public mempools.
On the other hand, a private mempool receives transactions from a portion of the web3 users, while a public mempool receives transactions from all web3 users. Thus, a disabled private mempool creates limited damage by victimizing the portion of users connected to the pool. The damage of a disable public mempool can be amplified by broadcasting the payload transactions across the network and reaching the public mempools of all nodes.
9 Conclusion
This paper presents mpfuzz, the first mempool fuzzer to find asymmetric DoS bugs by symbolically exploring mempool states and optimistically estimating the promisingness an intermediate state is in reaching bug oracles. Running mpfuzz on six major Ethereum clients leads to the discovery of new mempool vulnerabilities, which exhibit a wide variety of sophisticated patterns including stealthy mempool eviction and mempool locking.